OpenStudy (anonymous):

♥I will medal and fan♥ Which line shows the first error in the solution? A. line 2 B. line 3 C. line 4 D. line 5

OpenStudy (confxsedaf):

whats the solution?

OpenStudy (anonymous):

oh sorry hold on

OpenStudy (anonymous):

6(x+2)=24 (1) 6x+12=24 (2) 6x=12 (3) x=12 times 6 (4) x=72 (5)

OpenStudy (confxsedaf):

well if you look at number 4 they multiply 6x=12 which is wrong you have to divide. number 1, 2,3 look good but line 4 is wrong

OpenStudy (anonymous):

okay thank you I have 4 more is that okay

OpenStudy (confxsedaf):

sure

OpenStudy (anonymous):

What is the correct solution to the equation? 8(x + 2) = 24 A. 1 B. 3 C. 5 D. 64

OpenStudy (anonymous):

is it b

OpenStudy (anonymous):

no I mean a

OpenStudy (confxsedaf):

1.) 8(x + 2) = 24 2.) 8x+16=24 (distributive the 8 meaning multiply 8 by x and 8 by2) 3.) 8x=8 divide 4.) x=1

OpenStudy (confxsedaf):

sorry i forgot to subtract 16 by 16 and 16 by 24 8x=8

OpenStudy (anonymous):

oh okay thank you I get it

OpenStudy (confxsedaf):

yep. if u need help i'm here :)